OpenStudy (anonymous):

(2x + 4)(x - 6)? Could someone explain to me how to do this?

OpenStudy (anonymous):

FOIL method

OpenStudy (anonymous):

u know what that is?

OpenStudy (anonymous):

Are you simplifying it?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

i can tell you the answer

OpenStudy (anonymous):

(2x +4)(x-6) Multiply x and 2 Multiply x and 1 The x just gets copied along. The answer is x x 2*x evaluates to 2x 2*x+4 evaluates to 2x+4 x-6 evaluates to x-6 Multiplying 2x+4 by x-6 is a classic Algebra problem. Here, you are trying to multiply two binomials together (two expressions that each contain two terms). Your book might call this finding the "Product of Two Binomials". To work this problem, we'll use the "F.O.I.L." method. F.O.I.L. stands for First, Outer, Inner, Last. First, we'll multiply the two First terms, the 2x and x together. Multiply 2x and x Multiply the x and x Multiply x and x Combine the x and x by adding the exponents, and keeping the x, to get The answer is 2x × x = Second, we'll multiply the two Outer terms, the 2x and -6 together. Multiply 2x and -6 Multiply x and 1 The x just gets copied along. The answer is x x 2x × -6 = -12x Third, we'll multiply the two Inner terms, the 4 and x together. Multiply 4 and x Multiply 1 and x The x just gets copied along. x 4 × x = 4x -12x combines with 4x to give -8x Lastly, we'll multiply the two Last terms, the 4 and -6 together. Multiply 4 and -6 1 4 × -6 = -24 (2*x+4)*(x-6) evaluates to The final answer is

OpenStudy (anonymous):

2x^2 -8x - 24

OpenStudy (anonymous):

Thank you both for the help!